FOURIER TRANSFORM MICROWAVE SPECTROSCOPY OF ARGON CHLOROCYCLOBUTANE

JODI M. SZARKO, \underline{RANGANATHAN SUBRAMANIAN}, WALLACE C. PRINGLE, and STEW-ART E. NOVICK, Department of Chemistry, Wesleyan University, Middletown, CT 06459.

The microwave spectrum of chlorocyclobutane was reported in 1966^{a}. We have measured the microwave spectrum of argon chlorocyclobutane as part of our on-going research on bonding and dynamics of four-membered ring containing van der Waals complexes. Preliminary results show that the four-membered ring remains puckered in the complex. An analysis based partially on the nuclear quadrupole coupling projections of the ^{35}Cl nucleus will be presented.

aH. Kim and W. D. Gwinn J. Chem. Phys. 44, 865, (1966)